write the equation of a ellipse given the center (3,1) the foci (6,1) and the vertex (8,1_
Answer by lwsshak3(11628)

write the equation of a ellipse given the center (3,1) the foci (6,1) and the vertex (8,1)
**
This is an equation of an ellipse with horizontal major axis.
Its standard form: (x-h)^2/a^2+(y-k)^2/b^2=1, a>b, (h,k)=(x,y) coordinates of center
For given ellipse:
center: (3,1) (given)
a=5 (from 3 to 8)
a^2=25
..
c=3 (from 3 to 6)
c^2=9
..
c^2=a^2-b^2
b^2=a^2-c^2=25-9=16
b=4
..
Equation of given ellipse:
(x-3)^2/25+(y-1)^2/16=1